OpenStudy (anonymous):

A circle has radius 10 cm. Use an approximation for to find (a) the circumference of the circle and (b) the area of the circle to the nearest integer.

OpenStudy (texaschic101):

well...the circumference of a circle is : C = 2 * (pi) * r. so an approximate answer is : C = 2 * 3 * 10 = 60 area of a circle is : A = (pi)(r^2) so the approximate answer is : A = 3 * 100 = 300 I realize that pi is 3.14, but an approximate answer is not the exact answer.